[Hallc_running] [New Logentry] RC daily update

mnycz at jlab.org mnycz at jlab.org
Thu Aug 13 12:35:01 EDT 2020

Logentry Text:
It has been an eventful day shift
The Moller measurement took all of day and most of swing yesterday. Bill and Murchhana were able to finish up late swing. Thanks to Bill and Murchhana for staying late to finish!
The Moller result --> https://logbooks.jlab.org/entry/3826006

After Moller, we started back to production on the polarized Helium-3 cell

We have some had target related issues / hiccups
It was noticed at the beginning of owl that Laser Spectrum Monitor Peak disappeared -->  https://logbooks.jlab.org/entry/3826153
Melanie has spent the morning with the target experts trying to debug the QWP rotation stage controller

We tried to turn on the camera to see the QWP, but with no luck (it was already on. Brad suggested it may have gotten unplugged). We followed the instructions in --> https://hallcweb.jlab.org/wiki/index.php/Pol_He-3_Rotation_Stages#Power_to_camera_and_lamp_in_Optics_Box

 We took an escorted access and I found the power strip the camera and light were connected was turned off. After turning back on, the camera was up an running (for a few minutes at least)
As of 12:10 pm, Work is still being done on the QWP rotation stage controller to ensure we have control over it.

The NMR measurement decreased during OWL shift. 

We will meet today at 4:15pm via Bluejeans 861439920 : https://hallcweb.jlab.org/wiki/index.php/RC_Update_and_Run_Plan_for_Aug_13,_2020#Current_Run_Plan


This is a plain text email for clients that cannot display HTML.  The full logentry can be found online at https://logbooks.jlab.org/entry/3826544
-------------- next part --------------
An HTML attachment was scrubbed...
URL: <https://mailman.jlab.org/pipermail/hallc_running/attachments/20200813/ec658c5b/attachment.html>

More information about the Hallc_running mailing list